OpenStudy (anonymous):

9^x=1/5sqrt(3) Solve for x PLEASE HELP

OpenStudy (anonymous):

\[9^{x}=\frac{ 1 }{ \sqrt[5]{3} }\]

OpenStudy (jhannybean):

\[9^x = (3^2)^x = 3^{2x}\]\[\frac{1}{\sqrt[5]{3}} = \frac{1}{3^{1/5}} = 3^{-1/5}\]\[9^{x}=\frac{ 1 }{ \sqrt[5]{3} }\]\[3^{2x} = 3^{-1/5}\]Because the bases are the same, we can evaluate their powers.\[2x = -\frac{1}{5}\] Can you solve for x from here?

OpenStudy (anonymous):

I'm really bad at fractions i dont know how to divide that

OpenStudy (jhannybean):

Ok, \[\large -\frac{\frac{1}{5}}{2} \implies -\frac{1}{5} \cdot \frac{1}{2}\] when dividing a fraction by a number, you treat the fraction on top as the value of 1, so you can intuitively read (fraction)/number as (1)/number. Then you separate them and multiply the fraction ontop by the number on bottom (which you turned into a fraction in your head) Get it? :D

OpenStudy (anonymous):

so -1/10?

OpenStudy (jhannybean):

Yup :)